What is the discount yield, bond equivalent yield, and effective annual return on a $1 million T-bill that currently sells at 97.375 percent of its face value and is 75 days from maturity? (3 decimal places) (Q1)
Answer #1

Discount yield = (($1,000,000 − $973,750)/$1,000,000)(360/75) = 12.600%

Bond equivalent yield = (($1,000,000 − $973,750)/$973,750)(365/75) = 13.119%

Effective annual return = (1 + 0.13119/(365/75))365/75 − 1 = 13.821%